OpenStudy (anonymous):

Jim invested his savings in two investment funds. The amount he invested in Fund A was times as much as the amount he invested in Fund B. Fund A returned a profit and Fund B returned a profit. How much did he invest in Fund B, if the total profit from the two funds together was ?

OpenStudy (jdoe0001):

"The amount he invested in Fund A was times as much" ^ ? "Fund A returned a profit and Fund B returned a profit." ^ ^ ? ?

OpenStudy (anonymous):

so what is it

OpenStudy (jdoe0001):

dunno, the posting has a few blanks

OpenStudy (anonymous):

its beginning Algebra

OpenStudy (jdoe0001):

well, you do not have any amounts in the posting

OpenStudy (anonymous):

Jim invested his savings in two investment funds. The amount he invested in Fund A was 4 times as much as the amount he invested in Fund B. Fund A returned a 3% profit and Fund B returned a 4% profit. How much did he invest in Fund B, if the total profit from the two funds together was $1120?

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%\ of \ b=\cfrac{{\color{red}{ a}}}{100}\times b\implies (0.0a)b\) thus 3% of A will be 0.03A and 4% of B will be 0.04B "A was 4 times as much as the amount he invested in Fund B" so whatever B was, A was 4*B or 4B, thus A = 4B the total profit those 2 funds made was $1120 so one can say that \(\bf 0.03A+0.04B=1120\qquad {\color{blue}{ A=4B}}\qquad thus \\ \quad \\ 0.03{\color{blue}{ A}}+0.04B=112\implies 0.03({\color{blue}{4B}})+0.04B=1120\) solve for B

OpenStudy (anonymous):

so what the answer and how did u get it

OpenStudy (jdoe0001):

well, solve for B

OpenStudy (anonymous):

k i got that do u mind helping me with something else